# Exact fluctuation relation for open systems beyond the Jarzynski equality

### v1

@@ Abstract
Frozen coupling regime. The scenario of frozen coupling deserves particular attention. In most free energy difference calculations — both theoretical and experimental — the system remains in fixed coupling (the coupling strength $C$ between $\mathcal{S}$ and $\mathcal{E}$ is kept constant throughout the process) with its environment while the control parameter $\lambda(t)$ is driven. This is precisely the regime underlying the standard derivation of the JE, although the JE itself is agnostic to whether the protocol acts as a driving or a coupling. Within this common, physically motivated freezing scenario, the formal structure of our derivation simplifies considerably.
@@
For a fixed coupling protocol $C(t)\equiv C(0)$, the HMF increment along a trajectory reduces exactly to the difference of bare–system Hamiltonians,
@@
Equation (1) follows directly from Eq. (3), for constant $C$, the environmental contribution inside the logarithmic term cancels between the two endpoints. Substituting Eq. (1) into the trajectory equalities (16)–(17) yields
@@
Corollary 2 (Heat–work decomposition for frozen coupling).
We now discuss the thermodynamic structure hidden in Eqs. (1) and  (1). In the frozen coupling regimes, energetic changes in $\mathcal{S}$ arise from (i) external manipulation of $\lambda(t)$ and (ii) exchange with the environment through the microscopic evolution of $X_{\mathcal{E}}$. To inspect these contributions, we start from Eq. (1) and apply a simple algebraic insertion–subtraction of the initial energy at fixed protocol, $\mathcal{H}_{\mathcal{S}}\bigl(X_{\mathcal{S}}(0|X_{0}),\lambda(0)\bigr)$, which yields the following representation. For $C(t)\equiv C(0)$,
$\mathcal{H}_{\beta}^{*}\bigl(\mathcal{T}_{t_{\rm eq}}^{\mathcal{S}}(X_{0}),\lambda(t_{\rm eq}),C(0)\bigr)-\mathcal{H}_{\beta}^{*}\bigl(\mathcal{T}_{t_{\rm eq}}^{\mathcal{S}}(X_{0}),\lambda(0),C(0)\bigr)$
@@
The first term describes the energy exchange with the environment along the actual phase–space trajectory and defines the instantaneous heat rate, $\dot{Q}_{\mathcal{S}}$, whereas the second term accounts for the parametric energy input due to the protocol and defines the instantaneous work rate, $\dot{W}_{\mathcal{S}}$. Following the standard conventions of stochastic energetics [50], their time integrals give
@@
Here the gradient is evaluated with respect to $\mathcal{H}_{\mathcal{S}}$ frozen at $\lambda(0)$, whereas the $\dot{X}_{\mathcal{S}}$ belongs to the driven trajectory generated by $\lambda(t)$. Consequently, $II$ is not a thermodynamic heat unless $\dot{\lambda}=0$ for all $t$; it is a reference functional that projects the driven generalized velocity ($\dot{X}_{\mathcal{S}}$) onto the generalized force field ($\nabla_{X_{\mathcal{S}}}\mathcal{H}_{\mathcal{S}}$) of the initial stationary protocol. Because the reference and realized dynamics belong to distinct protocol layers, $II$ acts as a feedback-like correction without a definite thermodynamic notion. Combining Eqs. (25) and (26) with Eq. (2), the total bare–system increment becomes
@@
Substituting Eq. (27) into Eqs. (1)–(1) yields the compact heat–work representation of the endpoint equalities,
@@
GJE connection. The GJE restores $e^{-\beta\Delta F}$ under feedback by augmenting work with mutual information, providing an operational measure of information as a thermodynamic resource [21]. Our identities achieve restoration without invoking measurement–feedback a priori and without assuming LDB. The pathwise correction $Q_{\mathcal{S}}-II$ compensates dynamical asymmetry during driving, and the ensemble overlap $1+\chi^{2}(P^{\mathcal{S}}_{t_{\rm eq}}\parallel P^{\mathcal{S}}_{0})$ compensates endpoint mismatch. Formally, our equality is broader in scope, it holds without feedback and without LDB, while its structure is fully compatible with the GJE.
Figure 2: (A) System free energy difference $\Delta F_{\mathcal{S}}^{*}$ versus the final control value $\lambda(t_{\rm eq})$: circles (JE), squares (Eq. 1), triangles (Eq. 1), solid line (exact). Ramp/post-ramp protocol: $\lambda(t)$ increases linearly from $\lambda(0)$ to the target over $t_{\lambda}$, then is held fixed. (B) Representative run for $\lambda(0)=1\!\to\!\lambda(t_{\rm eq})=4$. Center: $\lambda(t)$. Top/bottom: ensemble trajectories of $x(t)$ and $p_{x}(t)$ (dashed line marks $t_{\lambda}$). Left/right: canonical marginals $P^{\mathcal{S}}_{0}(x,p_{x})$ and $P^{\mathcal{S}}_{t_{\rm eq}}(x,p_{x})$. During the ramp, dynamics are non-Liouvillian with damping $(\gamma_{\rm nl}^{\mathcal{S}},\gamma_{\rm nl}^{\mathcal{E}})$, post-ramp relaxation is governed by the underdamped Langevin equations, whose solution converges to the canonical distribution of the composite system, thereby satisfying the asymptotic equilibration postulate, Eq. (15).
@@
This equality, which corresponds to [Eq. (23) in [30]], contains no reference to the dynamics of the composite system. All relations derived in this work—independent of coupling strength or dynamical assumptions—reduce to this equilibrium partition function ratio. The standard JE can be viewed as the specific realization of this identity under Hamiltonian evolution of the composite system. For completeness, the detailed derivation linking Eq. (31) to the conventional JE is provided in Appendix, where we recast Jarzynski’s original argument using the present notation.
@@
Validation. To validate the framework, we examine how the proposed equalities hold beyond the regime where the JE applies. The model construction is inspired by the two–stage thought experiment introduced by Jarzynski in his reply [30] to Cohen’s critique [22]. In his argument, the control parameter is first driven from $\lambda(0)$ to $\lambda(t_{\lambda})$ (stage 1) and then held fixed at $\lambda=\lambda(t_{\lambda})$ for relaxation (stage 2). Two observers — one stopping at $t_{\lambda}$ (out of equilibrium) and the other at $t_{\mathrm{eq}}$ (after relaxation) — record the same work, since no work is performed once the driving protocol is fixed, and calculate the same $\Delta F_{\mathcal{S}}^{*}$, defined as the equilibrium free energy difference between the canonical states associated with $\lambda(0)$ and $\lambda(t_{\lambda})=\lambda(t_{\mathrm{eq}})$ via the HMF partition functions. As Jarzynski emphasized [30], ”whether or not we choose to include a relaxation stage has no bearing on the validity” of the JE. While this statement is entirely correct, it implicitly depends on a key assumption about the underlying dynamics. To probe this assumption directly, we construct a validation model consisting of a non–Liouvillian ramp (stage 1) that breaks phase–space volume preservation, followed by an underdamped Langevin relaxation (stage 2) that enforces the required equilibrium endpoint. In this scenario, the JE no longer holds because phase–space contraction during the ramp breaks the Liouvillian structure of the composite dynamics, preventing the equilibrium partition function ratio in Eq. (31) from reproducing the JE. For a detailed discussion of the validation model and protocol, see the Appendix.
@@
Conclusions: In this Letter, we establish exact fluctuation relations for open systems. Using the HMF, the free energy difference is expressed via exponential moments of the HMF shift and an endpoint chi-squared divergence (Thm. 1). The trajectory counterparts of that relation is obtained under the assumption of asymptotic equilibration (Thm. 2). In the frozen coupling regime the relations reduce to bare-system expressions (Cor. 1) with a clear heat–work decomposition (Cor. 2), and they include the JE and its feedback extensions (GJE) as limiting cases. Validation on a composite model with a non-Liouvillian ramp followed by underdamped Langevin relaxation shows that our endpoint and trajectory relations reproduce the exact free energy differences, demonstrating accuracy and practical applicability beyond the JE regime. The generality of this framework invites direct application to complex nonequilibrium settings where traditional FTs are fragile or inapplicable, including active matter [51, 52], biological processes [53], complex molecular and chemical environments [54], and open quantum systems [55]. In these regimes, our equalities provide a practical route to reconstruct free energy landscapes from realistic driving protocols without imposing idealized dynamical models or weak coupling assumptions. When applied to situations in which the JE is valid, our relations reduce to an exact re-expression of the JE that makes explicit the roles of endpoint overlap and system–environment interaction through the chi-squared factor and the HMF structure. In this sense, the JE becomes a diagnostic tool: deviations can be traced to insufficient overlap [56], rare-event sampling issues [57], or neglected coupling effects [50]. These diagnostics offer concrete guidance for designing and refining experimental and simulation protocols [58], and we expect them to support more reliable free energy estimation across a broad class of strongly driven open systems.
@@ Acknowledgements.
We gratefully acknowledge financial support by the DFG under Germany’s Excellence Strategy EXC 2089/1-390776260 (e-conversion).

### v2

@@ Abstract
The passage from Eqs. (15)-(16) to Eqs. (21)-(22) follows by inserting the trajectory representation of $P^{\mathcal{S}}_{t_{\mathrm{eq}}}$ and exchanging limits and integrals under the asymptotic equilibration condition (see [PRE]).
Frozen coupling regime. The scenario of frozen coupling deserves particular attention. In a broad class of nonequilibrium work protocols — both theoretical and experimental — the system remains in fixed coupling (the interaction strength $C$ between $\mathcal{S}$ and $\mathcal{E}$ is kept constant throughout the process) while the control parameter $\lambda(t)$ is driven. This is precisely the regime underlying the standard derivation of the JE [jarzynski1997equilibrium, jarzynski1997nonequilibrium, jarzynski2004nonequilibrium], although the JE itself is agnostic to whether the protocol acts through $\lambda$ or through $C$. Within this common, physically motivated freezing scenario, the formal structure of our derivation simplifies considerably. The complementary frozen-driving viewpoint, together with its direct connection to free energy perturbation (FEP), is developed in detail in the companion article [PRE].
@@
For a fixed coupling protocol $C(t)\equiv C(0)$, the HMF increment along a trajectory reduces exactly to the difference of bare-system Hamiltonians,
@@
Equation (1) follows directly from Eq. (Exact fluctuation relation for open systems beyond the Jarzynski equality), for constant $C$, the environmental contribution inside the logarithmic term cancels between the two endpoints. Substituting Eq. (1) into the trajectory equalities (21)-(22) yields
@@
Corollary 2 (Heat-work-feedback decomposition for frozen coupling).
We now discuss the thermodynamic structure hidden in Eqs. (1) and  (1). In the frozen coupling regimes, energetic changes in $\mathcal{S}$ arise from (i) external manipulation of $\lambda(t)$ and (ii) exchange with the environment through the microscopic evolution of $X_{\mathcal{E}}$. To inspect these contributions, we start from Eq. (1) and apply a simple algebraic insertion-subtraction of the initial energy at fixed protocol, $\mathcal{H}_{\mathcal{S}}\bigl(X_{\mathcal{S}}(0|X_{0}),\lambda(0)\bigr)$, which yields the following representation. For $C(t)\equiv C(0)$,
$\mathcal{H}_{\mathcal{S}}\bigl(\mathcal{T}_{t_{\rm eq}}^{\mathcal{S}}(X_{0}),\lambda(t_{\rm eq})\bigr)-\mathcal{H}_{\mathcal{S}}\bigl(\mathcal{T}_{t_{\rm eq}}^{\mathcal{S}}(X_{0}),\lambda(0)\bigr)$
@@
The first term describes the energy exchange with the environment along the actual phase-space trajectory and defines the instantaneous heat rate, $\dot{Q}_{\mathcal{S}}$, whereas the second term accounts for the parametric energy input due to the protocol and defines the instantaneous work rate, $\dot{W}_{\mathcal{S}}$. Following the standard conventions of stochastic energetics [talkner2020colloquium], their time integrals give
@@
Here the gradient is evaluated with respect to $\mathcal{H}_{\mathcal{S}}$ frozen at $\lambda(0)$, whereas the $\dot{X}_{\mathcal{S}}$ belongs to the driven trajectory generated by $\lambda(t)$. Consequently, $II$ is a reference functional that projects the driven generalized velocity ($\dot{X}_{\mathcal{S}}$) onto the generalized force field ($\nabla_{X_{\mathcal{S}}}\mathcal{H}_{\mathcal{S}}$) of the initial stationary protocol. Because the reference and realized dynamics belong to distinct protocol layers, $II$ acts as a feedback-like correction [sagawa2010generalized] without a definite thermodynamic notion. Combining Eqs. (30) and (31) with Eq. (2), the total bare-system increment becomes
@@
Substituting Eq. (32) into Eqs. (1)-(1) yields the heat-work-feedback representation of the endpoint equalities,
@@
GJE connection. The GJE restores $e^{-\beta\Delta F}$ under feedback by augmenting work with mutual information, providing an operational measure of information as a thermodynamic resource [sagawa2010generalized]. Our identities achieve restoration without invoking measurement-feedback a priori and without assuming LDB. The pathwise correction $Q_{\mathcal{S}}-II$ compensates dynamical asymmetry during driving, and the ensemble overlap $1+\chi^{2}(P^{\mathcal{S}}_{t_{\rm eq}}\parallel P^{\mathcal{S}}_{0})$ compensates endpoint mismatch. Formally, our equality is broader in scope, it holds without feedback and without LDB, while its structure is fully compatible with the GJE [sagawa2010generalized].
Figure 1: (A) System free energy difference $\Delta F_{\mathcal{S}}^{*}$ versus the final control value $\lambda(t_{\rm eq})$: circles (JE), squares (Eq. (1)), triangles (Eq. (1)), solid line (exact). Ramp/post-ramp protocol: $\lambda(t)$ increases linearly from $\lambda(0)$ to the target over $t_{\lambda}$, then is held fixed. (B) Representative run for $\lambda(0)=1\!\to\!\lambda(t_{\rm eq})=4$. Center: $\lambda(t)$. Top/bottom: ensemble trajectories of $x(t)$ and $p_{x}(t)$ (dashed line marks $t_{\lambda}$). Left/right: canonical marginals $P^{\mathcal{S}}_{0}(x,p_{x})$ and $P^{\mathcal{S}}_{t_{\rm eq}}(x,p_{x})$. During the ramp, dynamics are non-Liouvillian with damping $(\gamma_{\rm nl}^{\mathcal{S}},\gamma_{\rm nl}^{\mathcal{E}})$, post-ramp relaxation is governed by the underdamped Langevin equations, whose solution converges to the canonical distribution of the composite system, thereby satisfying the asymptotic equilibration postulate, Eq. (19).
@@
This equality, which corresponds to [Eq. (23) in [jarzynski2004nonequilibrium]], contains no reference to the dynamics of the composite system. All relations derived in this work—independent of coupling strength or dynamical assumptions—reduce to this equilibrium partition function ratio. The standard JE can be viewed as the specific realization of this identity under Hamiltonian evolution of the composite system [jarzynski2004nonequilibrium]. For completeness, the detailed derivation linking Eq. (36) to the conventional JE is provided in the Appendix, where we recast Jarzynski’s original argument using the present notation.
@@
Validation. To validate the framework, we examine how the proposed equalities hold beyond the regime where the JE applies. The model construction is inspired by the two-stage thought experiment introduced by Jarzynski in his reply [jarzynski2004nonequilibrium] to Cohen’s critique [cohen2004note]. In his argument, the control parameter is first driven from $\lambda(0)$ to $\lambda(t_{\lambda})$ (stage 1) and then held fixed at $\lambda=\lambda(t_{\lambda})$ for relaxation (stage 2). Two observers — one stopping at $t_{\lambda}$ (out of equilibrium) and the other at $t_{\mathrm{eq}}$ (after relaxation) — record the same work, since no work is performed once the driving protocol is fixed, and calculate the same $\Delta F_{\mathcal{S}}^{*}$, defined as the equilibrium free energy difference between the canonical states associated with $\lambda(0)$ and $\lambda(t_{\lambda})=\lambda(t_{\mathrm{eq}})$ via the HMF partition functions. As Jarzynski emphasized [jarzynski2004nonequilibrium], ”whether or not we choose to include a relaxation stage has no bearing on the validity” of the JE. While this statement is entirely correct, it implicitly depends on a key assumption about the underlying dynamics. To probe this assumption directly, we construct a validation model consisting of a non-Liouvillian ramp (stage 1) that breaks phase-space volume preservation, followed by an underdamped Langevin relaxation (stage 2) that enforces the required equilibrium endpoint. In this scenario, the JE no longer holds because phase-space contraction during the ramp breaks the Liouvillian structure of the composite dynamics, preventing the equilibrium partition function ratio in Eq. (36) from reproducing the JE. By contrast, the same partition function ratio leads directly to the endpoint relations in Thm. 1, without any dynamical assumptions, and the trajectory form in Thm. 2 follows once asymptotic equilibration is imposed as a boundary condition. For a detailed discussion of the validation model and protocol, see the Appendix.
@@
Conclusions: In this Letter, we establish exact fluctuation relations for open systems. Using the HMF, the free energy difference is expressed via exponential moments of the HMF shift and an endpoint chi-squared divergence (Thm. 1). The trajectory counterparts of that relation are obtained under the assumption of asymptotic equilibration (Thm. 2). In the frozen coupling regime the relations reduce to bare-system expressions (Cor. 1) with a clear heat-work-feedback decomposition (Cor. 2), and they include the JE and its feedback extensions (GJE) as limiting cases. Validation on a composite model with a non-Liouvillian ramp followed by underdamped Langevin relaxation shows that our endpoint and trajectory relations reproduce the exact free energy differences, demonstrating accuracy and practical applicability beyond the JE regime. The generality of this framework invites direct application to complex nonequilibrium settings where traditional FTs are fragile or inapplicable, including active matter [johnsrud2025fluctuation, bebon2025thermodynamics], biological processes [woodside2014reconstructing], complex molecular and chemical environments [leighton2025flow], and open quantum systems [beyer2025operational]. In these regimes, our equalities provide a practical route to reconstruct free energy landscapes from realistic driving protocols without imposing idealized dynamical models or weak coupling assumptions. When applied to situations in which the JE is valid, our relations provide an exact re-expression of the JE that makes explicit the roles of endpoint overlap and system-environment interaction. In this sense, our relations can be interpreted as an enriched form of the JE, providing diagnostic tools that clarify how practical failures of the JE can arise, for example due to insufficient endpoint overlap [schmidt2025free], rare-event sampling limitations [kuang2020gamma], or unaccounted strong-coupling effects [talkner2020colloquium]. Together, this framework can potentially offer concrete guidance for designing and refining experimental and simulation protocols [casert2024learning], for controlling dissipation and optimizing driving in soft and active matter [johnsrud2025fluctuation, bebon2025thermodynamics, engel2023optimal], for improving nonequilibrium binding free energy calculations and molecular design workflows [kuang2020gamma, baumann2021challenges, serra2025nonequilibrium, serra2025path], and for advancing the operation and characterization of strongly coupled quantum devices and engines [cangemi2024quantum].
@@ Acknowledgements.
We gratefully acknowledge financial support by the DFG under Germany’s Excellence Strategy EXC 2089/2-390776260 (e-conversion).
End Matter
Appendix: We collect detailed proofs of the theorems and corollaries stated in the main text. Each result is reproduced in full mathematical form to ensure completeness and clarity of the derivations.
Figure 2: (A) Schematic of the composite $\mathcal{S}+\mathcal{E}$ at a fixed bath temperature $T=\beta^{-1}/k_{B}$. The control $\lambda(t)$ (schematic purple spring) and the coupling $C(t)$ (schematic black spring) are driven independently and become constant at times $t_{\lambda}$ and $t_{C}$, respectively; the composite system then relaxes to equilibrium at $t_{\rm eq}$. The system free energy difference $\Delta F_{\mathcal{S}}^{*}(\beta)$ refers to the canonical endpoints $(\lambda(0),C(0))\!\to\!(\lambda(t_{\rm eq}),C(t_{\rm eq}))$ and is defined through the HMF partition functions. (B) An initial microstate $X_{0}\!\sim\!P_{0}(X,\beta)$ on $\Gamma(0)$ evolves under the kinematic map $X(t|X_{0})=\mathcal{T}_{t}(X_{0})$ through the ramp and the post-ramp relaxation, yielding the asymptotic density $P_{t_{\rm eq}}(X,\beta)$ on $\Gamma(t_{\rm eq})$. The frozen coupling case, $C(t)\equiv C(0)$, is the regime used in our validation.
Figure 3: Same plotting conventions as the Fig. 1. During the ramp we set $\gamma_{\rm nl}^{\mathcal{S}}=\gamma_{\rm nl}^{\mathcal{E}}=0$.
Table 1: Numerical inputs for the validation runs (frozen coupling; non-Liouvillian ramp, under damped Langevin relaxation). All quantities are in reduced, dimensionless units.
Quantity Value
Mass ($m$) $1.0$
System–environment coupling $C$ $1.0$
$\beta$ $1.0$
System–environment coupling $C$ $1.0$
Environment frequency $\omega$ $1.0$
Initial protocol $\lambda(0)$ $1.0$
Ramp duration $t_{\mathrm{ramp}}$ $10.0$
Relaxation duration $t_{\mathrm{eq}}-t_{\lambda}$ $20.0$
Ramp time step $\Delta t$ $2\times 10^{-3}$
Relaxation time step $\Delta t_{\mathrm{relax}}$ $2\times 10^{-2}$
Non-Liouvillian drag (system) $\gamma_{\mathrm{nl}}^{\mathcal{S}}$ $20.0$
Non-Liouvillian drag (environment) $\gamma_{\mathrm{nl}}^{\mathcal{E}}$ $5.0$
Langevin damping (relaxation) $\gamma$ $2.0$
Initial ensemble size $N_{\mathrm{traj}}$ $30000$
Relax integrator BAOAB [kieninger2022gromacs]
JE as limiting case. Starting from the Eq. (36), and assuming the composite system evolves under Hamilton’s equations generated by $\mathcal{H}_{\mathcal{S}+\mathcal{E}}$. The instantaneous energy change along a trajectory $X(t|X_{0})$ obeys
$\frac{\mathrm{d}}{\mathrm{d}t}\,\mathcal{H}_{\mathcal{S}+\mathcal{E}}(X(t|X_{0}),\lambda(t),C(t))$
$=\underbrace{\nabla_{X}\mathcal{H}_{\mathcal{S}+\mathcal{E}}\cdot\dot{X}}_{\text{advection}}+\frac{\partial\mathcal{H}_{\mathcal{S}}}{\partial\lambda}\,\dot{\lambda}+\frac{\partial\mathcal{V}_{\mathcal{S}\mathcal{E}}}{\partial C}\,\dot{C}.$ (37)
The advection term vanishes because the Hamiltonian flow is symplectic, $\nabla_{X}\mathcal{H}_{\mathcal{S}+\mathcal{E}}\cdot\dot{X}=\{\mathcal{H}_{\mathcal{S}+\mathcal{E}},\mathcal{H}_{\mathcal{S}+\mathcal{E}}\}=0$. Integrating over $t\in[0,t_{\mathrm{eq}}]$ gives [jarzynski2004nonequilibrium]
$\mathcal{H}_{\mathcal{S}+\mathcal{E}}(X(t|X_{0}),\lambda(t),C(t))-\mathcal{H}_{\mathcal{S}+\mathcal{E}}(X(0|X_{0}),\lambda(0),C(0))$
$=\int_{0}^{t_{\mathrm{eq}}}{\rm d}t\,\Big(\frac{\partial\mathcal{H}_{\mathcal{S}}}{\partial\lambda}\dot{\lambda}+\frac{\partial\mathcal{V}_{\mathcal{S}\mathcal{E}}}{\partial C}\dot{C}\Big)=:W_{\lambda}+W_{C}.$ (38)
Using the above equation and Liouville’s theorem, which guarantees a unit Jacobian for the change of variables ($X_{0}\mapsto X$), Jarzynski [jarzynski2004nonequilibrium] demonstrated that the equilibrium partition function ratio for the composite system can be expressed as
$\Big\langle e^{-\beta(W_{\lambda}+W_{C})}\Big\rangle_{X_{0},\beta}=\frac{Z_{\mathcal{S}+\mathcal{E}}(\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}),\beta)}{Z_{\mathcal{S}+\mathcal{E}}(\lambda(0),C(0),\beta)}.$ (39)
For frozen coupling $C(t)\equiv C(0)$, Eq. (39) reduces to
$e^{-\beta\Delta F_{\mathcal{S}}^{*}}=\Big\langle e^{-\beta W_{\lambda}}\Big\rangle_{X_{0},\beta},$ (40)
which is the conventional JE expressed in the HMF notation. $\square$
Validation model and protocol: Model specification. We validate Eqs. (1) and (1) on an analytically tractable composite system consisting of a system with double well potential [whitelam2025improving, hanggi1990reaction, li2021equilibrium, saito1976relaxation, sun2003equilibrium] coupled to a harmonic environment. The system Hamiltonian is $\mathcal{H}_{\mathcal{S}}(x,p_{x};\lambda)=\frac{1}{2m}p_{x}^{2}+U_{\mathcal{S}}(x;\lambda)$, where $U_{\mathcal{S}}(x;\lambda)=\tfrac{1}{4}(x^{2}-\lambda)^{2}$, the environment Hamiltonian $\mathcal{H}_{\mathcal{E}}(y,p_{y})=\frac{1}{2m}p_{y}^{2}+\tfrac{1}{2}\omega^{2}y^{2}$, and the interaction $\mathcal{V}_{\mathcal{S}\mathcal{E}}(x,y;C)=Cxy$ with coupling constant $C$. The total Hamiltonian follows Eq. (Exact fluctuation relation for open systems beyond the Jarzynski equality). Eq. (Exact fluctuation relation for open systems beyond the Jarzynski equality) gives $\mathcal{H}_{\beta}^{*}(x,p_{x};\lambda,C)$, the corresponding marginal and partition function follow from $P^{\mathcal{S}}(X_{\mathcal{S}},\beta)={e^{-\beta\,\mathcal{H}_{\beta}^{*}(X_{\mathcal{S}},\lambda,C)}}/{\mathcal{Z}_{\mathcal{S}}^{*}(\lambda,C,\beta)}$ and $\mathcal{Z}_{\mathcal{S}}^{*}(\lambda,C,\beta)=\int\mathrm{d}X_{\mathcal{S}}\,e^{-\beta\,\mathcal{H}_{\beta}^{*}(X_{\mathcal{S}},\lambda,C)}$. The composite system is initialized in the canonical ensemble at $(\lambda(0),C(0))$, ensuring exact sampling from Eq. (11) (see Fig. 2). Two-stage protocol. We work in the frozen-coupling regime $C(t)\equiv C(0)$ and vary only $\lambda(t)$ (see Fig. 2). During the ramp $0\leq t\leq t_{\lambda}$, $\lambda(t)$ changes linearly from $\lambda(0)$ to $\lambda(t_{\lambda})$ (see Fig. 1). The composite evolves under non-Liouvillian deterministic dynamics $\dot{x}=p_{x}/m,\;\dot{p}_{x}=-\partial_{x}U_{\mathcal{S}}(x;\lambda(t))-Cy-\gamma_{\mathrm{nl}}^{\mathcal{S}}p_{x}/m,\;\dot{y}=p_{y}/m,\;\dot{p}_{y}=-\omega^{2}y-Cx-\gamma_{\mathrm{nl}}^{\mathcal{E}}p_{y}/m$, where $\gamma_{\mathrm{nl}}^{\mathcal{S}}$ and $\gamma_{\mathrm{nl}}^{\mathcal{E}}$ denote constant damping coefficients used only during the ramp. At $t=t_{\lambda}$, $\lambda(t)$ is held fixed and the composite relaxes under underdamped Langevin dynamics [hanggi1990reaction] obeying the FDT: $\dot{x}=p_{x}/m,\;\dot{p}_{x}=-\partial_{x}U_{\mathcal{S}}(x;\lambda(t_{\mathrm{eq}}))-Cy-\gamma p_{x}/m+\eta_{x}(t),\;\dot{y}=p_{y}/m,\;\dot{p}_{y}=-\omega^{2}y-Cx-\gamma p_{y}/m+\eta_{y}(t)$, where $\gamma$ is the friction coefficient [zwanzig2001nonequilibrium]. The terms $\eta_{x}(t)$ and $\eta_{y}(t)$ are independent standard Gaussian white noises with, $\langle\eta_{i}(t)\rangle=0$ and $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=(2\gamma/\beta)\delta_{ij}\delta(t-t^{\prime})$ for $i,j\in\{x,y\}$. These equations converge to the canonical distribution associated with the frozen Hamiltonian $\mathcal{H}_{\mathcal{S}+\mathcal{E}}(x,p_{x},y,p_{y};\lambda(t_{\mathrm{eq}}),C)$, satisfying the asymptotic equilibration postulate Eq. (19). Estimators. After relaxation, the relations Eqs. (1) and (1) are evaluated. The ”Exact” curve is the equilibrium free energy difference from Eq. (35). For the JE estimator we accumulate the ramp work
$W_{\lambda}=\int_{0}^{t_{\lambda}}\partial_{\lambda}U_{\mathcal{S}}(x_{t};\lambda_{t})\,\dot{\lambda}\,\mathrm{d}t,$ (41)
with frozen $C$, and compute, $e^{-\beta\Delta F_{\mathcal{S}}^{*}}=\langle e^{-\beta W_{\lambda}}\rangle$. Fig. 1 reports $\Delta F_{\mathcal{S}}^{*}$ as a function of $\lambda(t_{\mathrm{eq}})$: the JE estimator deviates because the Liouvillian assumption is violated, whereas the endpoint equalities reproduce the exact equilibrium free energies. As a consistency check, when we set $\gamma_{\rm nl}^{\mathcal{S}}=\gamma_{\rm nl}^{\mathcal{E}}=0$ (Hamiltonian/Liouvillian ramp), the JE work estimator, the endpoint equalities, and the equilibrium reference all coincide, recovering the standard JE (see Fig.3). For completeness, all numerical parameters used for the simulations are listed in Table. 1.